# README: Bridging Medical Jargon and Lay Understanding for Patient Education through Data-Centric NLP

### v1

@@ 4 Method
4.1 Dataset Construction
In this study, we utilized a collection of EHRs sourced from an anonymized institution. The identification of medical jargon within these EHRs was performed using the MedJEx, as detailed in Kwon et al. (2022). Domain experts provided lay definitions for these jargon terms, forming the dataset known as README. For training our models, it was necessary to define these jargon terms in a general yet scientifically accurate manner. To this end, we employed UMLS to generate general definitions. UMLS is a set of files and software that brings together many health and biomedical vocabularies and standards to enable interoperability between computer systems. We used the Scispacy library to access UMLS definitions of jargon terms. we used the en_core_sci_lg model to obtain the data. we tried other models in Scispacy but found all the other models to produce the same result as the en_core_sci_lg model  [^2] https://github.com/allenai/scispacy.
@@ 4.1 Dataset Construction
Given the complexity of medical terminology, it’s noteworthy that some terms in UMLS are associated with multiple definitions. This reflects the reality that the meaning and application of a medical term can vary based on its contextual use. To accurately select the most appropriate general definition for each context, we utilized the SentenceBERT similarity score, as proposed by Reimers and Gurevych (2019), to identify the most fitting definition.
@@
Our solution to address these scenarios involved a unified approach, given the impracticality of tackling each case individually. As illustrated in Figure 3, we extracted the UMLS definitions for all phrases within a jargon term. This process yielded a list of strings, providing a comprehensive general definition for each term. There might not be a UMLS definition for all the words in the jargon term. In that case, we are concatenating the word directly. Subsequently, we applied LLMs for the necessary data-cleaning steps.
@@
The initial README dataset has a lot of unusable data. There are many reasons why a data point might not be useful. The main reasons are (i) the lay definition provided by the annotator is not correct or missing. (ii) The jargon term is missing. (iii)The general definition obtained from UMLS is too scientific or not relevant to the jargon term. Using this data for training would be bad for the training as these data points will make the model worse. So these data points are not considered. This preliminary cleaning brings out data from 350K to 308K.
4.2 Examiner-Augmenter-Examiner (EAE)
4.2.1 Examiner (expert-annotated data)
Initially, basic data cleaning, as outlined in Section 4.1, was applied. To enhance this, we employed GPT-3.5-turbo, using a few-shot learning approach with seven examples - four demonstrating acceptable data points and three showing unacceptable ones. These prompts served as the ‘Human’ element in our Human-AI-in-the-loop model, as depicted in Figure 2 and detailed in Appendix Algorithm 1. The prompts are detailed in Appendix Table 5. We choose GPT-3.5-turbo here because recent work Remy and Demeester (2023) and our evaluation (4.2.4 and 5.3) show that the definitions it generates for medical terms can reach a human-satisfying level. Post-cleaning, approximately 40% of UMLS general definitions were deemed suitable for model training. The suitable data points were archived in README-v2, while the unsuitable ones were stored in README-v3. To minimize computing costs, we excluded EHRs and took all the unique data points. This was done because, in many cases, we have the same jargon term, general definition from UMLS, and lay definition but different EHR’s. This reduced our dataset from 308,239 to 51,623 data points. Upon processing these through GPT-3.5-turbo, 11,765 were classified as ‘good’ quality general definitions, and 39,856 as ‘bad’. A subsequent SQL join with contextual data removed earlier resulted in 113,659 ‘good’ and 194,580 ‘bad’ general definitions. After eliminating duplicates, the final count for bad general definitions in README-v3 was 177,140.
README-EAE
       inputs : README-v1
       output : Final README-v6 dataset with best general definitions and lay definitions
       $Initialize:$ $exam_{p}=examinerpromptforgpt3$;
       $aug_{p}=augmenterpromptsforChatGPT$;
       foreach $datapoint:README-v1$ do
            if $exam_{p}(datapoint)==yes$ then
                   README-v2.add(datapoint);
             end if
            else
                   README-v3.add(datapoint);
             end if
       end foreach
      foreach $datapoint:README-v3$ do
            $temp=aug_{p}(datapoint)$;
             README-v4.add(temp);
             if $exam_{p}(temp)==yes$ then
                   README-v5.add(temp);
             end if
            else
                   README-v6.add(temp);
             end if
       end foreach
      return README-v6;
Algorithm 1 Algorithm for Data Cleaning
4.2.2 Augmenter
Given the low yield of usable UMLS definitions, we employed ChatGPT, a recent LLM by OpenAI, to augment our dataset. The augmentation process, a critical part of our Data-centric pipeline, involved the system prompt: "Generate a general definition of the term for a professional medical student." Accompanied by two examples(Table 5), this step aimed to create complex yet general definitions suitable for professional medical students. The outcome of this process was README-v4, containing 171,831 newly generated definitions.
4.2.3 Examiner (AI-synthetic data)
The ChatGPT-generated definitions underwent a second cleaning round using the same methodology as in Examiner (expert-annotated data). Here, approximately 56% of the ChatGPT definitions were found suitable for model training, with the remaining being either contextually inappropriate or incompatible with the expert-provided lay definitions. The final tally was 96,585 ‘good’ and 75,068 ‘bad’ general definitions, stored in README-v5 and README-v6, respectively.
4.2.4 Data Quality

### v2

@@ 3 Method
3.1 README Data Collection
The dataset source is a collection of publicly available deidentified EHR notes from hospitals affiliated with an anonymized institution. Herein, 18,178 sentences were randomly sampled, and domain experts then annotated the sentences for medical jargon and corresponding lay definition.
3.2 Lay Definition Annotation
Domain-experts read each sentence and identified as medical jargon terms that would be considered difficult to comprehend for anyone no greater than a 7th-grade education  [^3] The rule of thumb is that if a candidate term has a lay definition comprehensible to a 4-7th grader as judged by Flesch-Kincaid Grade Level (Solnyshkina et al., 2017), the candidate term is included as a jargon term.. Overall, 51,623 unique (medical jargon term, lay definition) pairs with 308,242 mentions  [^4] So there are 308,242 data points with (EHR context, medical jargon term, lay definition) format. in the EHR repository have been annotated by complying with the annotation guidelines presented in Appendix A.
3.3 General Definition Retrieval
We then employed the Scispacy library  [^5] We used the Scispacy en_core_sci_lg model to obtain the data. https://github.com/allenai/scispacy to retrieve the corresponding UMLS definitions Lindberg et al. (1993) of these annotated medical jargon terms as the general definitions. We follow the retrieval and preprocessing steps in Appendix C to filter the valid general definitions for the README dataset. This preliminary cleaning results in 308,242 data points in the README-exp (e.g., expert annotated) dataset, each consisting of a clinical note context, a medical jargon term, a corresponding lay definition, and a corresponding general definition, as shown in Figure 2.
3.4 Examiner-Augmenter-Examiner (EAE)
3.4.1 Examiner (expert-annotated data)
Initially, basic data cleaning, as outlined in Appendix C, was applied. To enhance this, we employed GPT-3.5-turbo, using a few-shot learning approach with seven examples - four demonstrating acceptable data points and three showing unacceptable ones. These prompts served as the ‘Human’ element in our Human-AI-in-the-loop model, as depicted in Figure 2 and detailed in Algorithm 1. The prompts are detailed in Table 6. We choose GPT-3.5-turbo here because our evaluation (Section 4.3 and Appendix D) shows that the definitions it generates for medical terms can reach a human-satisfying level. Post-cleaning  [^6] More GPT Running Details in Examiner step can be found on Appendix B.4, approximately 39% of UMLS general definitions were deemed suitable by the Examiner (e.g., GPT-3.5-turbo). The suitable 113,659 data points were archived in R-exp_good, while the unsuitable 177,140 ones were stored in R-exp_bad.
3.4.2 Augmenter
Given the low yield of usable UMLS definitions, we employed GPT-3.5-turbo to augment our dataset. The augmentation process, a critical part of our Data-centric pipeline, involved the system prompt: “Generate a general definition of the term.” This step, accompanied by two examples (Table 6), aimed to create correct general definitions but may not be suitable for laypeople (similar to the UMLS definitions). The outcome of this process was R-syn (e.g., AI synthetic), containing 171,831 newly generated definitions.
3.4.3 Examiner (AI-synthetic data)
The ChatGPT-generated definitions underwent a second cleaning round using the same methodology as in Section 3.4.1 Examiner (expert-annotated data). Here, approximately 56% of the ChatGPT definitions were found suitable for model training, with the remaining being either contextually inappropriate or incompatible with the expert-provided lay definitions. The final tally was 96,668 ‘good’ and 75,175 ‘bad’ general definitions, stored in R-syn_good and R-syn_bad, respectively.
After the end of the EAE, we sampled 500 medical jargon terms each from the R-exp_good and R-syn_good datasets. We conducted human verification  [^7] The details about Data Quality Checking and Train/Eval/Test Split can be found in Appendix D on corresponding data points for these 1000 jargon terms. We obtained a high human agreement for both the R-exp_good and R-syn_good datasets’ quality. After correcting individual invalid data, we used this portion of the data as evaluation and test data (in a 1:1 ratio). We used the remaining R-exp_good and R-syn_good datasets as training data, ensuring that these 1,000 medical jargon terms would not appear in the training data. Table 5 shows the overall dataset statistics of all the README versions across the pipeline.
3.5 Integration of Synthetic and Expert Data
We adopted four distinct sampling strategies to integrate the AI-synthetic training data (e.g., R-syn_good) into the expert-annotated training data (e.g., R-exp_good):
  • RANDOM: This approach randomly selected N entries from the R-syn_good dataset. This is the baseline for our subsequent three heuristic methods.
  • SEMANTIC: For semantic-based sampling, we utilized SentenceTransformers  [^8] We used default model all-MiniLM-L6-v2 in https://github.com/UKPLab/sentence-transformers.. Renowned for its text semantic analysis efficiency, this model enabled us to measure the semantic similarity between lay definitions in R-exp_good and R-syn_good datasets. We ranked synthetic data based on these scores, considering higher scores as indicative of greater semantic closeness to expert annotations.
  • MODEL: In model-based sampling, we used models initially trained on the R-exp_good dataset to generate definitions for the R-syn_good dataset. We employed the ROUGE_L F1 score to evaluate the alignment between model-generated and actual R-syn_good lay definitions. This technique aids in mitigating training challenges associated with data heterogeneity. It enriches the dataset with examples that enhance the model’s convergence towards the desired distribution (e.g., expert-annotated lay definitions).
4 Experiments
4.1 Automatic Evaluation Metrics
We evaluate the efficacy of our model in producing lay definitions by contrasting them with the ground-truth reference lay definitions, utilizing the ROUGE Lin (2004) and METEOR Banerjee and Lavie (2005) metrics. However, they are based on the exact word overlap and therefore provide insight into the informativeness of the generated lay definitions but do not necessarily reflect their factual accuracy Maynez et al. (2020). We also employ Scispacy to extract medical concepts from the model-generated and the reference lay definitions. We then compute the F1 Scores for these concept lists, referred to as UMLS-F1, to specifically measure the factuality of the generated content  [^9] The details about UMLS-F1 can be found in Appendix E.
4.2 Experimental Setting
We use the following base models: GPT-2, DistilGPT2, BioGPT, and Llama2 in our experiments  [^10] More details about Experimental Setting can be found in Appendix F.. We use the following symbols:
  1. 1.
  2. 2.
  3. 3.
  4. 4.
We use one jargon, ’EDG’, as an example and show the input prompt of different settings J2L, J+C2L, J+G2L, J+C+G2L on Table 7.
Our experiments were divided into five distinct parts. In Set-1, we aimed to evaluate the performance gap on the J2L task between open-source base models and GPT3.5/4 models. We do one-shot prompting for Llama2  [^11] Because smaller open source language models DistilGPT2/GPT2/BioGPT do not have the ability to complete instruction following under zero-shot or few-shot settings, we only compare the results of Llama2 as a representative of open source language models with GPT3.5/4 to see the gap., GPT-3.5-turbo, and GPT-4 with prompts in Table 6. Set-2 explored the varying data quality across different versions within our EAE pipeline, where we fine-tuned the base models on J+G2L task. Set-3 focused on evaluating the effects of various data selection strategies on data augmentation outcomes. To do this, we ranked AI-synthetic data (R-syn_good) using different methods in Section 3.5, selecting either the top N entries with the highest scores (e.g., ‘SEMANTIC’ in Table 2) or the bottom N entries with the lowest scores (e.g., ‘SEMANTIC_r’). These selections were then evenly combined with expert-annotated data (R-exp_good) at a one-to-one ratio. Following this, we fine-tuned the base models using these diverse, mixed datasets to determine the impact of each selection method on model performance. In Set-4, we investigated the effects of incorporating different types of information (EHR context and UMLS-retrieved general definition) into the model inputs. In Set-5, we fine-tuned models of varying sizes (ranging from DistilGPT2-88M to Llama2-7B) and compared the outcomes with GPT-3.5-turbo using best settings learned from previous Set-2, Set-3, and Set-4.
4.3 Results
Figure 3: One-shot performances on jargon2lay.
We start with our experiments with some base models’ performance on the J2L dataset with Set-1. In Appendix D, we showed a high level of human evaluators’ agreement for the definitions generated by GPT-3.5-turbo, which underlines its efficacy in crafting human-readable explanations for biomedical concepts. Consequently, GPT-3.5-turbo serves as a strong baseline for quality in README lay definition generation, indicative of a standard that meets human satisfaction. Despite this, our analysis of Llama2, illustrated in Figure 3, reveals a significant performance gap. This discrepancy underscores the critical necessity for enhancing the capabilities of Llama2 and other open-source models to achieve high-quality output in lay definition generation tasks, thereby improving patient education together with systems like NoteAid. Additionally, we observed that GPT-3.5-turbo and GPT4 exhibited comparable proficiency in this task.
4.3.1 Effectiveness of EAE pipeline
J+G2L ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1 Rank
R-exp 23.94 8.95 22.79 17.75 12.83 4
R-exp_good 26.99 10.76 25.57 21.19 16.88 2
R-exp_bad 22.25 7.46 21.09 17.18 11.72 5
R-syn_good 25.71 9.75 24.41 20.23 16.01 3
R-exp+syn_good 29.82 13.14 28.47 24.42 20.09 1
Table 1: Various README versions data performance.
In Set-2, we focus on the efficacy of different data versions when finetuning with the GPT-2 model. The results, as reflected in Table 1, indicate that high-quality expert data (R-exp_good) demonstrates clear superiority over unexamined expert data (R-exp), emphasizing the crucial role of ECE-examiner in enhancing data quality. Furthermore, high-quality synthetic data (R-syn_good) outperforms the unexamined expert data (R-exp), underscoring the significant value of ECE-augmenter. Notably, the combination of R-exp_good and R-syn_good shows improved performance over R-exp_good alone, suggesting that including synthetic data is beneficial. This composite approach of R-exp_good+R-syn_good leading the rank underscores the efficacy of our EAE pipeline.
4.3.2 Expert and Synthetic Data Integration
ROUGE1 J2L J+C2L J+G2L J+C+G2L
RANDOM 19.03 19.65 26.21 26.97
SYNTAX(R) 19.82(-2.28) 20.65(-2.78) 27.56(-8.36) 28.33(-8.84)
SEMANTIC(R) 19.74(-1.11) 19.94(-1.39) 26.42(-0.92) 27.19(-1.96)
MODEL(R) 20.18(-0.8) 20.46(-1.88) 27.9(-2.37) 28.65(-4.22)
UMLS-f J2L J+C2L J+G2L J+C+G2L
RANDOM 8.05 9.09 15.78 16.72
SYNTAX(R) 8.53(-0.11) 9.99(-1.00) 17.76(-5.11) 17.42(-5.54)
SEMANTIC(R) 8.98(-0.38) 9.07(-0.69) 15.93(-0.02) 16.74(-0.93)
MODEL(R) 9.18(-0.63) 9.4(-0.78) 18.34(-2.5) 18.45(-2.55)
Table 2: Different data selection methods performance. The values in parentheses represent the difference between the corresponding X_r (bottom N) and X (top N), i.e., X_r - X. The smaller this value, the more it indicates that the X method can select higher quality synthetic data for data augmentation; conversely, the closer this value is to 0, the more it suggests that the X method cannot identify the higher-quality synthetic data for training.
In Set-3, we explored the effects of various data selection strategies on data augmentation outcomes. We found the results of SEMANTIC are closer to RANDOM than SYNTAX and MODEL, and SEMANTIC_r - SEMANTIC is close to 0. However, significant differences can be observed in SYNTAX and SYNTAX_r, MODEL and MODEL_r. Table 2 highlights two main findings: firstly, data selection is crucial, as all methods have better performance with higher-ranked data (e.g., SYNTAX, SEMANTIC, MODEL) over lower (e.g., SYNTAX_r, SEMANTIC_r, MODEL_r). All three methods, SYNTAX, SEMANTIC, and MODEL, have better results than the RANDOM baseline. Secondly, SYNTAX and MODEL are more effective in selecting higher-quality synthetic data for data augmentation than SEMANTIC and RANDOM. We selected SYNTAX as the default data augmentation method for subsequent experiments.
4.3.3 Retrieval-augmented Generation
ROUGE1 ROUGE2 ROUGEL METEOR UMLS-F1
Without data augmentation (R-exp_good)
J2L 19.47 6.06 18.53 14.74 8.76
J+C2L 19.40 6.40 18.38 15.12 9.24
J+G2L 26.99 10.76 25.57 21.19 16.88
J+C+G2L 27.58 11.31 26.35 21.72 17.12
Data augmentation (R-exp+syn_good with SYNTAX)
J2L 21.98 7.42 20.88 16.98 10.95
J+C2L 22.13 7.90 21.04 17.54 10.71
J+G2L 29.82 13.14 28.47 24.42 20.09
J+C+G2L 29.89 13.48 28.49 24.65 20.27
Table 3: Efficacy of incorporating EHR context and general definition in input data. The retrieved general definitions significantly aid the overall performance of the model (ROUGE and METEOR) and also reduce hallucinations (UMLS factuality score).
Set-4 results underscore the significant improvement in model performance when input data is enriched with UMLS-retrieved general definitions. As illustrated in Table 3, regardless of whether we utilize only expert-annotated data or data augmentation with AI-synthetic data, including general definitions consistently enhances effectiveness. This finding confirms the value of RAG with the general definition in the lay definition generation task. Meanwhile, adding EHR context to the input data yields a moderate impact on model performance.
4.3.4 Model Performances Against ChatGPT
In Set-5, we observed that LLAMA2-7B’s ROUGE-1 and UMLS-F1 metrics surpassed GPT-3.5-turbo in the J2L task post-training. For the J+G2L setting, DistilGPT2-88M demonstrated equivalent results to GPT-3.5-turbo, while BioGPT’s performance exceeded it, and LLAMA2-7B significantly outperformed the GPT-3.5-turbo. These findings, as depicted in Figure 4, emphasize the effectiveness of open-source, mobile-adapted small models when appropriately fine-tuned with high-quality datasets, offering a promising avenue for deploying lightweight yet powerful NLP tools in mobile healthcare applications to help patient education.
5 Human Evaluation
5.1 Human Evaluation settings
Our human evaluation was conducted by 5 human evaluators  [^12] Since the generated lay definition is provided to lay people, five people without medical background were found to conduct the human evaluation here.. We randomly selected 50 pairs of (jargon, generated lay definitions) from the test dataset for this human evaluation. The task for evaluators was to reference the expert definitions and choose a binary preference among the following four groups of definitions: 1) DistilGPT2-J2L vs. GPT-3.5-turbo, 2) DistilGPT2-J+C+G2L vs. GPT-3.5-turbo, 3) LLAMA2-J2L vs. GPT-3.5-turbo, 4) LLAMA2-J+C+G2L vs. GPT-3.5-turbo. After we get judgments from multiple people per instance, we do not aggregate their labels before calculating the win rate but count them individually.
5.2 Human Evaluation Results
Figure 5: Human evaluation results (win rate %).
As shown in Figure 5, although the result of adding EHR context and general definition (DistilGPT2-J+C+G2L) is better than DistilGPT2-J2L, the win rate of the two DistilGPT2 models’ output is still significantly behind the result of GPT-3.5-turbo. For LLAMA-2, generating lay definitions directly from jargon is still not as good as GPT-3.5-turbo, but adding context and general definitions is of great help. Human evaluators prefer LLAMA2-J+C+G2L more than GPT-3.5-turbo. There are some inconsistencies between the results of human evaluation and automatic evaluation. We further interviewed our medical experts about the reasons for our system win or loss cases and concluded the following conclusions to help future improvements: 1. While all our in-house systems perform satisfactorily, GPT-3.5-turbo stands out for its flexibility and user-friendliness. It excels at elaborating complex medical terms, offering detailed explanations and practical examples to improve comprehension. 2. Recent advancements Cai et al. (2023); Zhang et al. (2023b) reveal ChatGPT’s role in enhancing patient education through interactive formats like NoteAid-interactive Zhang et al. (2023a). It enables patients to actively ask questions and seek clarifications while the AI tailors responses to aid their understanding. This interactive approach, absent in traditional dictionary-style definitions like our README dataset, calls for next-step model distillation work or further refinement in aligning the in-house system’s outputs with patient preferences. 3. Additionally, developing automatic metrics aligning closely with human evaluation is another critical next step.
6 Related Work
The evolution of patient-centric healthcare necessitates simplified patient access to medical information. Tools like NoteAid and MedJEx have initiated efforts to make EHR content more comprehensible (Chen et al., 2018; Kwon et al., 2022). However, sentence-level text simplification efforts have been expanded to larger datasets that capture broader biomedical contexts (Jonnalagadda et al., 2010; Guo et al., 2021; Devaraj et al., 2021). Recent research has similarly focused on the development of datasets and methods for text simplification, primarily at the sentence level, with the expansion into datasets capturing broader contexts of biomedical abstracts (Cao et al., 2020; Lu et al., 2023; Goldsack et al., 2022; Luo et al., 2022b). Notably, efforts such as the CELLS, PLABA, and AGCT datasets have contributed significantly to this domain, providing extensive resources for training models capable of translating scientific discourse into lay language (Guo et al., 2024; Attal et al., 2023; Remy and Demeester, 2023). Our work diverges from these existing efforts by introducing the README dataset, an expansive collection explicitly designed for context-aware lay definitions, addressing the nuanced task of generating patient-friendly definitions directly from medical terms, filling a critical gap in patient education resources.
In line with advancing the quality of generated texts, we have embraced Retrieval-Augmented Generation (RAG) to mitigate common issues in natural language generation, such as "hallucinations" (Karpukhin et al., 2020; Shuster et al., 2021). Two main categories of information retrieval methods have been used to augment the generation of biomedical natural language generation, definition-based and embedding-based retrieval techniques (Guo et al., 2024; Alambo et al., 2022; Moradi and Ghadiri, 2018; Xiong et al., 2024). Our RAG belongs to the definition-based retrieval technique.
Finally, our contributions distinctly highlight the integration of a robust, data-centric Human-AI pipeline that improves data quality and the efficiency of models trained on the README dataset. This innovative pipeline leverages the Data-centric AI framework, navigating through phases of collection, labeling, preparation, reduction, and augmentation to build a dataset that is both expansive and representative Zha et al. (2023); Ng et al. (2021); Whang et al. (2023). The process begins with meticulous data collection and expert labeling, ensuring a foundation of high-quality, domain-specific data (Section 3.1 - 3.3). During the preparation phase, raw data undergoes rigorous cleaning and transformation, readying it for effective model training Krishnan and Wu (2019). The dataset is then enriched through strategic data augmentation techniques, incorporating verified quality AI-synthetic data to expand its scope and utility (Section 3.4). Furthermore, data reduction strategies are employed to select the most suitable instances for integration, enhancing the dataset’s overall effectiveness (Section 3.5). Through these meticulous stages, the README dataset not only supports but significantly enhances the capabilities of smaller, open-source models, allowing them to match or even exceed the performance of larger proprietary models like ChatGPT in specific healthcare applications.
7 Conclusions
Our study underscores the potential of NLP to democratize medical knowledge, enabling patient-centric care by simplifying complex medical terminology. Developing the README dataset and implementing a data-centric pipeline has improved dataset quality and expanded AI training possibilities. Our experiments show that small open-source models can match advanced close-source LLMs like ChatGPT with well-curated data. README will be open to the community as an important lay dictionary for patient education. We hope our work can help the innovative patient education community advance toward a future where all patients can easily understand their health information.
8 Limitations and Ethical Considerations
Appendix A Annotation Guideline
The dataset was annotated for medical jargon and lay definition by six domain experts from medicine, nursing, biostatistics, biochemistry, and biomedical literature curation  [^13] The annotator agreement scores can be found in Appendix B.. Herein, the annotators applied the following rules for identifying what was jargon and how to write a suitable lay definition:
Rule 1. Medical terms that would not be recognized by about 4 to 7th graders, or that have a different meaning in the medical context than in the lay context (homonym) were labeled. For example:
  • accommodate: When the eye changes focus from far to near.
  • antagonize: A drug or substance that stops the action or effect of another substance.
  • resident: A doctor who has finished medical school and is receiving more training.
  • formed: Stool that is solid.
Rule 2. Terms that are not strictly medical, but are frequently used in medicine. For example:
  • "aberrant", "acute", "ammonia", "tender", "intact", "negative", "evidence"
Rule 3. When jargon words are commonly used together, or together they mean something distinct or are difficult to understand from the individual parts quickly were labeled. For example:
  • vascular surgery: Medical specialty that performs surgery on blood vessels.
  • airway protection: Inserting a tube into the windpipe to keep it wide open and prevent vomit or other material from getting into the lungs.
  • posterior capsule: The thin layer of tissue behind the lens of the eye. It can become cloudy and blur vision.
  • right heart: The side of the heart that pumps blood from the body into the lungs.
  • intracerebral hemorrhage: A stroke.
Rule 4. Terms whose definitions are widely known (e.g., by a 3rd grader) do NOT need to be labeled. For example:
  • “muscle”, “heart”, “pain”, “rib”, “hospital”
Rule 4.1 When in doubt, label the term. For example:
  • “colon”, “immune system”
Appendix B Evaluation of the Annotation
An observational study was performed to evaluate the annotators’ reliability in identifying jargon and providing lay definitions, and assess the agreement of the dataset annotators with each other and laypeople.
B.1 Data Collection and Setting
For evaluation, twenty sentences were randomly selected from deidentified inpatient EHR notes in the EHR repository of one hospital affiliated with an anonymized institution. Sentences consisting only of administrative data, sentences less than ten words long, and sentences substantially indistinguishable from another sentence were filtered out.
Note that the annotators had never seen the sampled sentences. The twenty sentences were made up of 904 words in total. Common words were discarded so as not to inflate the calculated agreement. These consisted of all pronouns, conjunctions, prepositions, numerals, articles, contractions, months, punctuation, and the most common 25 verbs, nouns, adverbs, and adjectives. Terms occurring more than one time in a sentence were counted only once. Furthermore, multi-word terms were counted as single terms to ameliorate the double-counting issue. Two members of the research team determined multi-word terms by consensus. In this work, multi-word terms were defined as adjacent words that represented a distinct medical entity (examples: “PR interval”, “internal capsule”, “acute intermittent porphyria”), were commonly used together (examples: “hemodynamically stable”, “status post”, “past medical history”) and terms that were modified by a minor word (examples: “trace perihepatic fluid”, “mild mitral regurgitation”, “rare positive cells”, “deep pelvis”). After applying these rules, 325 candidate medical jargon terms and their lay definition were utilized. The laypeople comprised 270 individuals recruited from Amazon Mechanical Turk (MTurk) (Aguinis et al., 2021).
B.2 Annotation Reliability
The results showed that there was good agreement among annotators (Fleiss’ kappa = 0.781). The annotators had high sensitivity (91.7%) and specificity (88.2%) in identifying jargon terms and providing suitable lay definitions as determined by the laypeople (the gold standard).
B.3 Details about Jargons and Lay Definitions Statistics in README-exp
B.4 GPT Running Details in Examiner step
To optimize computing resources, we streamlined our dataset by removing EHR context and focusing solely on unique data points with (medical jargon term, lay definition, general definition) format in this Examiner step  [^14] This was done because, in many cases, we have the same jargon term, general definition from UMLS, and lay definition but different EHR contexts. We can reduce the amount of GPT running if we ignore the EHR context difference.. This reduced our dataset from 308,242 to 51,623 data points. Upon processing these through Examiner, 11,765 were classified as ‘good’ quality general definitions, and 39,856 as ‘bad’. We subsequently performed an SQL join operation, integrating the ‘good’ and ‘bad’ datasets with the previously removed EHR context data, which resulted in 113,659 ‘good’ and 194,580 ‘bad’ data points with (EHR context, medical jargon terms, lay definition, general definition) format. After eliminating duplicates, the final count for bad data points in R-exp_bad was 177,140.
Appendix C General Definition Retrieval and Preprocessing
@@
UMLS Lindberg et al. (1993) is a set of files and software that combines many health and biomedical vocabularies and standards to enable interoperability between computer systems. Given the complexity of medical terminology, it’s noteworthy that some terms in UMLS are associated with multiple definitions. This reflects the reality that a medical term’s meaning can vary depending on its contextual use. To accurately select the most appropriate general definition for each context, we utilized the SentenceBERT Reimers and Gurevych (2019) similarity score between lay definitions and all possible UMLS definitions available for a jargon term to identify the most fitting definition.
@@
Our solution to address these scenarios involved a unified approach, given the impracticality of tackling each case individually. As illustrated in Figure 6, we extracted the UMLS definitions for all phrases within a jargon term. This process yielded a list of strings, providing a comprehensive general definition for each term. We then concatenate these strings in a comma-separated manner to get one string used as a general definition. There might not be a UMLS definition for all the words in the jargon term. In that case, we are concatenating the word directly.
@@
The initial README dataset has a lot of unusable data. There are many reasons why a data point might not be useful. The main reasons are (i) the lay definition provided by the annotator is not correct or missing. (ii) The jargon term is missing. (iii)The general definition obtained from UMLS is too scientific or not relevant to the jargon term. This was done by making sure there were no empty column values and removing parsing issues of comma-separated files. Using this data for training would be bad for the training as these data points will make the model worse. So these data points are not considered. This preliminary cleaning brings out data from 350K to 308K.
Appendix D Data Quality Checking and Train/Eval/Test Split after EAE Pipeline